OpenStudy (anonymous):

A certain elevator cab has a total run of 190 m and a maximum speed of 305 m/min, and it accelerates from rest and then back to rest at 1.22 m/s2. (a) How far does the cab move while accelerating to full speed from rest? (b) How long does it take to make the nonstop 190 m run, starting and ending at rest?

sam (.sam.):

Vo = 0 10.6m = 1/2at^2 ------------------------------------ or Vx = Vox + at 5.0833 = 0 + 1.22 t t = 4.17 s =========================== 2t to find the time to both accelerate and decelerate. Distance traveled at 5.0833 m/s will be d = 190m -2(10.6m) = 168.8 m Time to travel 168.8 m at 5.0833 m/s T = d/v = 168.8m / 5.0833 m/s = 33.21 s Total time is Total time = 33.21 +2( 4.17 ) = 41.5 s